The inverse demand curve for a monopolist’s product is P = -Q + 8. What is the maximum
revenue the monopolist can earn if it cannot price discriminate?
1
Expert's answer
2017-03-06T15:20:05-0500
TR=P*Q=Q(-Q+8)=-Q^2+8Q
Let’s assume Q=1: -1+8=7
Q=2: -4+16=12
Q=3: -9+24=15
Q=4: -16+32=16
Q=5: -25+40=15
Q=6: -36+48=12
The maximum revenue the monopolist can earn is 16